# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ Appendix D Deferred Proofs
$\left|\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|=\left|\frac{1}{T}\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|\leq\frac{\epsilon}{T}.$
Substituting $\epsilon=\bar{v}\sqrt{T\ln T}$, the above is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, and by also substituting this $\epsilon$ into (26) and by Lemma D.2, we have $\Pr(E)\geq\Pr(E_{1})-O(1/T^{2})$, giving the result. ∎
@@
Proof.
It was proven in [fikioris2023online] that $\beta\mapsto v_{i}^{\star}(\beta)$ is concave. The lemma statement follows from concavity and the fact that $v_{i}^{\star}(0)=0$. ∎
With those lemmas, we can now prove that each player following a $\beta_{i}$-aggressive strategy and give their utility guarantee. We use $p_{i}$ to denote player $i$’s interim allocation probability, which we define in Definition A.2. Equivalently, $p_{i}$ denotes $p_{i}(\nu)$ as defined in (24) when $\nu$ is the distribution over $S_{\neq i}$ where $j\in S_{\neq i}$ independently with probability $\beta_{j}$. The below theorem is Propositions A.1 and A.2 combined. It is a direct generalization of Proposition 4.1, which can be obtained by setting $\beta_{i}=\alpha_{i}$ for each $i$. When using $\beta_{i}=\alpha_{i}$ and Lemmas 6.2 and 6.3 to set the allocation probabilities, we also obtain Theorem 4.3.
Theorem D.5.
@@
Proof.
Suppose every agent $i$ is using a $\beta_{i}$-aggressive strategy. Let $E_{1}$ be the event that no agent runs out of budget. By Lemma D.2, $\Pr(E_{1})\geq 1-O(1/T^{2})$. Using Lemma D.3, there is an event $E$ of probability at least $1-O(1/T^{2})$ on which
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\geq p_{i}v^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$ (27)
Now we upper bound the utility of any deviating strategy by player $i$, still assuming players $j\neq i$ are following a $\beta_{j}$-aggressive strategy. Any strategy $(b_{i}^{t})$ used by player $i$ in the actual game satisfies the budget constraint $\sum_{t=1}^{T}b_{i}^{T}\leq\beta_{i}(1+\delta_{i}^{T})T$ almost surely. In particular, it satisfies the budget constraint in expectation, so we can use Lemma 5.1 to conclude that under the strategy $(b_{i}^{t})$ in the imaginary game,
@@
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]&=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}}]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq p_{i}v_{i}^{\star}(\beta_{i}(1+\delta_{i}^{T}))+O\left(\frac{1}{T^{2}}\right).\end{split}$ (28)
By (27) and (28), by deviating from a $\beta_{i}$-aggressive strategy, player $i$ can only gain an additive utility difference of
@@
By substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$ and using Lemma D.4, this implies that this additive difference is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, thus proving the theorem.
D.3 Deferred Proofs of Theorems B.1, B.2, and B.3
In all the proofs in this subsection, we use the following notation. Let $\tilde{b}_{i}^{t}$ be the bids of agent $i$ corresponding to a $\beta_{i}$-aggressive strategy in the imaginary game introduced in Section D.2 where there is no budget constraint that agree with the actual bids $b_{i}^{t}$ at times $t$ where agent $i$ has budget. Let $U_{i}[t,T]$ denote the utility of player $i$ gained at time $t$ under the policy $\tilde{b}_{i}^{t}$ if only the $n$ constraints $\sum_{t=1}^{T}b_{k}^{t}\leq\beta_{k}(1+\delta_{k}^{T})T$ were enforced. If agent $i$ deviates to a policy $(b_{i}^{t})$, we let $U_{i}^{\prime}[t]$ and $U_{i}^{\prime}[t,T]$ be analogous to $U_{i}[t]$ and $U_{i}[t,T]$ for the deviating policy.
Proof of Theorem B.1.
By Lemma D.2, the event $E_{1}^{s}$ that no agent runs out of budget at time $s$ has probability at least $1-O(1/s^{2})$. By the union bound, the event $E_{1}=\bigcap_{s=\lceil\sqrt{t}\rceil}^{t}E_{1}^{s}$ that no agent runs out of budget at any time $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-\sum_{s=\lceil\sqrt{t}\rceil}^{t}(1-\Pr(E_{1}^{s}))=1-O(1/\sqrt{t})$. By Lemma D.3, on an subevent $E$ of $E_{1}$ with probability at least $1-O(1/\sqrt{t})$,
@@
Suppose agent $i$ deviates to a policy $(b_{i}^{t})^{\prime}$. By Theorem D.5,
@@
By substituting $t=T$, we see that everyone playing a $\beta_{i}$-aggressive strategy is indeed an $O\left(\sqrt{\frac{\log T}{T}}\right)$-equilibrium. ∎
Proof of Theorem B.2.
Assume without loss of generality that the agents $j\neq i$ never bid when they’re out of budget. As in proof of Theorem B.1, the event $E_{1}$ that agent $i$ does not run out of budget at any time between $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-O(1/\sqrt{t})$. Then, on $E_{1}$, no one runs out of budget between time $\lceil\sqrt{t}\rceil$ and $t$. Using Theorem 4.3, there is an subevent $E$ of $E_{1}$ of probability at least $1-O(1/\sqrt{t})$ on which
@@
Proof of Theorem B.3.
By Lemma D.2 and the Borel-Cantelli Lemma, there exists a random time $t_{0}$ such that $\sum_{s=1}^{t}\tilde{b}_{i}^{s}\leq\beta_{i}(1+\delta_{i}^{t})t$ for all $t>t_{0}$ and all agents $i$ where $t_{0}<\infty$ almost surely. Suppose agent $i$ deviates to a policy $(b_{i}^{s})^{\prime}$. Using Theorem D.5,
@@
For the utility claim, let $\tilde{U}_{i}[s]$ denote the utility of agent $i$ in the imaginary game where no budget constraints are enforced and all agents $j$ are following a $\beta_{j}$-aggressive strategy. Observe that the $\tilde{U}_{i}[s]$ are i.i.d. $\mathrm{Bernoulli}(p_{i}v^{\star}(\beta_{i}))$. By the strong law of large numbers, $\frac{1}{t}\sum_{s=1}^{t}\tilde{U}_{i}[s]\overset{\mathrm{a.s}}{\to}p_{i}v^{\star}(\beta_{i})$. The utilities in the actual game satisfy
@@
which has the same limit as $t\to\infty$. ∎
D.4 Deferred Proof of Lemma 6.1
@@
Proof.
Suppose agent $i$ has arbitrary value distribution $\mathcal{F}_{i}$. Construct the policy $\pi_{i}$ as follows. At each time $t$, agent $i$ will sample $\hat{V}_{i}[t]\sim\mathrm{Bernoulli}((\rho_{i}^{\alpha_{i}})^{\star}(V_{i}[t]))$, where $(\rho_{i}^{\alpha_{i}})^{\star}$ is agent $i$’s $\alpha_{i}$-ideal utility probability function with value distribution $\mathcal{F}_{i}$. (We define $\alpha_{i}$-ideal utility probability function in Definition A.1; informally, it is $1$ if $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of $\mathcal{F}_{i}$ and 00 otherwise.) Then, agent $i$ will bid if and only if $\hat{\pi}_{i}$ would bid with the Bernoulli value $\hat{V}_{i}[t]$. In other words, the policy $\pi_{i}$ is simply following the policy $\hat{\pi}_{i}$ but with the Bernoulli values $\hat{V}_{i}[t]$ instead of the actual values.
@@
Therefore, $M[t]$ is a $\mathcal{G}_{t}$-martingale. Let $\bar{v}$ be an upper bound on the distribution $\mathcal{F}_{i}$ (recall we assumed value distributions are bounded). By the Azuma-Hoeffding inequality, for any $\epsilon>0$,
$\Pr(M[T]\leq-\epsilon)\leq\exp\left(-\frac{2\epsilon^{2}}{\bar{v}^{2}T}\right).$
Set $\epsilon=\bar{v}\sqrt{T\ln T}$, so that the above is at most $O(1/T^{2})$. Let $E_{2}$ be the event that the above does not occur and let $E=E_{1}\cap E_{2}$. We have $\Pr(E)\geq 1-O(1/T^{2})$. On $E$,
@@
thereby proving the lemma. ∎
Now we can prove Lemma 6.1.
Proof of Lemma 6.1.
@@
This will be at most $\alpha_{i}\bar{p}(1-\alpha_{i})T+O(\sqrt{T\log T})$ on an event $E_{1}$ of probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by Azuma-Hoeffding applied to $M_{i}^{T}$. Also, agent $i$ will bid at least $\alpha_{i}T-(\sqrt{T\log T})$ times on some event $E_{2}$ or probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by standard Chernoff bounds. Let $E=E_{1}\cap E_{2}$. We lower bound agent $i$’s utility on $E$. Appealing to Lemma D.6, we assume without loss of generality that agent $i$ has a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution. Then, the $\alpha_{i}$-aggressive strategy is simply the strategy of bidding when $V_{i}[t]=1$ if there is budget left. Agent $i$’s utility can then be bounded below on $E$ as
@@
Because the $\alpha_{i}$-ideal utility under a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution is $\alpha_{i}$, this establishes the lemma. ∎
D.5 Completion of the Proof of Lemma 6.3
@@
where the inequality comes from the inequalities $-2+2\prod_{j\notin I}(1-y_{j})\leq 0$ and $-1+X\leq 0$.
@@
$\begin{split}(y_{j_{1}}-y_{j_{2}})\left(\frac{\partial f}{\partial y_{j_{1}}}-\frac{\partial f}{\partial y_{j_{2}}}\right)=\frac{1}{2}\prod_{i\in I}(1-x_{i})\prod_{j\notin I\cup\{j_{1},j_{2}\}}(1-y_{j})\left(2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\right).\end{split}$ (29)
By Lemma D.7 applied to the function $x\mapsto\frac{x}{1-x}$ on $[0,1]$,
$2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\geq 2X-\frac{\sum_{i\in I}x_{i}}{1-\sum_{i\in I}x_{i}}(1-X)=2X-\frac{X}{1-X}(1-X)=X\geq 0$
Therefore, (29) is nonnegative so $f((x_{i})_{i\in I},\cdot)$ is Schur-convex.
D.6 Deferred Proofs of Propositions 6.4 and 6.5
Notice that Proposition 6.4 is a special case of Lemma 6.5, so we only need to prove Lemma 6.5.
Proof of Lemma 6.5.
@@
For any single agent $j$, their bids are i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$ across time conditioned on them having budget remaining. Let $E$ be the event that no agent $j\neq i$ runs out of budget. For the same reason as in Lemma D.2, the probability that agent $j$ runs out of budget is at most $O(1/T^{2})$, so $\Pr(E)\geq 1-O(1/T^{2})$. Using the idea and notation of the imaginary game where budgets are not enforced as introduced in Section D.2, agent $i$’s expected utility can be bounded as
@@
By the strategy of the other agents, (we use the notation for $\beta$-ideal utility as in Definition A.1)
@@
using Lemma D.1 for the first inequality and Lemma D.4 for the second. The above and (D.6) imply the lemma statement. ∎
D.7 Deferred Proof of Theorem 6.6
Proof of Theorem 6.6.
@@
Using the above and Lemma 6.5,

### v2

@@ E Deferred Proofs
$\left|\frac{1}{T}\sum_{t=1}^{T}U_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|=\left|\frac{1}{T}\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})\right|\leq\frac{\varepsilon}{T}.$
Substituting $\varepsilon=\bar{v}\sqrt{T\ln T}$, the above is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, and by also substituting this $\varepsilon$ into (E.5) and by Lemma˜E.2, we have $\Pr(E)\geq\Pr(E_{1})-O(1/T^{2})$, giving the result.
@@
Proof 16.
It was proven in [fikioris2023online] that $\beta\mapsto v_{i}^{\star}(\beta)$ is concave. The lemma statement follows from concavity and the fact that $v_{i}^{\star}(0)=0$.
With those lemmas, we can now prove that each player following a $\beta_{i}$-aggressive strategy and give their utility guarantee. We use $p_{i}$ to denote player $i$’s interim allocation probability, which we define in Definition˜B.3. Equivalently, $p_{i}$ denotes $p_{i}(\nu)$ as defined in (E.3) when $\nu$ is the distribution over $S_{\neq i}$ where $j\in S_{\neq i}$ independently with probability $\beta_{j}$. The below theorem is Propositions˜B.2 and B.4 combined. It is a direct generalization of Proposition˜3.3, which can be obtained by setting $\beta_{i}=\alpha_{i}$ for each $i$. When using $\beta_{i}=\alpha_{i}$ and Lemmas˜5.2 and 5.3 to set the allocation probabilities, we also obtain Theorem˜3.6.
Theorem E.5.
@@
Proof 17.
Suppose every agent $i$ is using a $\beta_{i}$-aggressive strategy. Let $E_{1}$ be the event that no agent runs out of budget. By Lemma˜E.2, $\Pr(E_{1})\geq 1-O(1/T^{2})$. Using Lemma˜E.3, there is an event $E$ of probability at least $1-O(1/T^{2})$ on which
@@
(E.6) $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]\geq p_{i}v^{\star}(\beta_{i})-O\left(\sqrt{\frac{\log T}{T}}\right).$
Now we upper bound the utility of any deviating strategy by player $i$, still assuming players $j\neq i$ are following a $\beta_{j}$-aggressive strategy. Any strategy $(b_{i}^{t})$ used by player $i$ in the actual game satisfies the budget constraint $\sum_{t=1}^{T}b_{i}^{T}\leq\beta_{i}(1+\delta_{i}^{T})T$ almost surely. In particular, it satisfies the budget constraint in expectation, so we can use Lemma˜4.1 to conclude that under the strategy $(b_{i}^{t})$ in the imaginary game,
@@
(E.7) $\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]]&=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}}]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[U_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\boldsymbol{1}_{E_{1}^{c}}]\\ &\leq p_{i}v_{i}^{\star}(\beta_{i}(1+\delta_{i}^{T}))+O\left(\frac{1}{T^{2}}\right).\end{split}$
By (E.6) and (E.7), by deviating from a $\beta_{i}$-aggressive strategy, player $i$ can only gain an additive utility difference of
@@
By substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$ and using Lemma˜E.4, this implies that this additive difference is at most $O\left(\sqrt{\frac{\log T}{T}}\right)$, thus proving the theorem.
E.4 Deferred Proofs from section˜C
In all the proofs in this subsection, we use the following notation. Let $\tilde{b}_{i}^{t}$ be the bids of agent $i$ corresponding to a $\beta_{i}$-aggressive strategy in the imaginary game introduced in subsection˜E.3 where there is no budget constraint that agree with the actual bids $b_{i}^{t}$ at times $t$ where agent $i$ has budget. Let $U_{i}[t,T]$ denote the utility of player $i$ gained at time $t$ under the policy $\tilde{b}_{i}^{t}$ if only the $n$ constraints $\sum_{t=1}^{T}b_{k}^{t}\leq\beta_{k}(1+\delta_{k}^{T})T$ were enforced. If agent $i$ deviates to a policy $(b_{i}^{t})$, we let $U_{i}^{\prime}[t]$ and $U_{i}^{\prime}[t,T]$ be analogous to $U_{i}[t]$ and $U_{i}[t,T]$ for the deviating policy.
Proof 18 (Proof of Theorem˜C.1).
By Lemma˜E.2, the event $E_{1}^{s}$ that no agent runs out of budget at time $s$ has probability at least $1-O(1/s^{2})$. By the union bound, the event $E_{1}=\bigcap_{s=\lceil\sqrt{t}\rceil}^{t}E_{1}^{s}$ that no agent runs out of budget at any time $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-\sum_{s=\lceil\sqrt{t}\rceil}^{t}(1-\Pr(E_{1}^{s}))=1-O(1/\sqrt{t})$. By Lemma˜E.3, on an subevent $E$ of $E_{1}$ with probability at least $1-O(1/\sqrt{t})$,
@@ E.4 Deferred Proofs from section˜C
Suppose agent $i$ deviates to a policy $(b_{i}^{t})^{\prime}$. By Theorem˜E.5,
@@
By substituting $t=T$, we see that everyone playing a $\beta_{i}$-aggressive strategy is indeed an $O\left(\sqrt{\frac{\log T}{T}}\right)$-equilibrium.
Proof 19 (Proof of Theorem˜C.2).
Assume without loss of generality that the agents $j\neq i$ never bid when they’re out of budget. As in proof of Theorem˜C.1, the event $E_{1}$ that agent $i$ does not run out of budget at any time between $\lceil\sqrt{t}\rceil$ and $t$ has probability at least $1-O(1/\sqrt{t})$. Then, on $E_{1}$, no one runs out of budget between time $\lceil\sqrt{t}\rceil$ and $t$. Using Theorem˜3.6, there is an subevent $E$ of $E_{1}$ of probability at least $1-O(1/\sqrt{t})$ on which
@@
Proof 20 (Proof of Theorem˜C.3).
By Lemma˜E.2 and the Borel-Cantelli Lemma, there exists a random time $t_{0}$ such that $\sum_{s=1}^{t}\tilde{b}_{i}^{s}\leq\beta_{i}(1+\delta_{i}^{t})t$ for all $t>t_{0}$ and all agents $i$ where $t_{0}<\infty$ almost surely. Suppose agent $i$ deviates to a policy $(b_{i}^{s})^{\prime}$. Using Theorem˜E.5,
@@
For the utility claim, let $\tilde{U}_{i}[s]$ denote the utility of agent $i$ in the imaginary game where no budget constraints are enforced and all agents $j$ are following a $\beta_{j}$-aggressive strategy. Observe that the $\tilde{U}_{i}[s]$ are i.i.d. $\mathrm{Bernoulli}(p_{i}v^{\star}(\beta_{i}))$. By the strong law of large numbers, $\frac{1}{t}\sum_{s=1}^{t}\tilde{U}_{i}[s]\overset{\mathrm{a.s.}}{\to}p_{i}v^{\star}(\beta_{i})$. The utilities in the actual game satisfy
@@
which has the same limit as $t\to\infty$.
E.5 Deferred Proof of Lemma˜5.1
@@
Proof 21.
Suppose agent $i$ has arbitrary value distribution $\mathcal{F}_{i}$. Construct the policy $\pi_{i}$ as follows. At each time $t$, agent $i$ will sample $\hat{V}_{i}[t]\sim\mathrm{Bernoulli}((\rho_{i}^{\alpha_{i}})^{\star}(V_{i}[t]))$, where $(\rho_{i}^{\alpha_{i}})^{\star}$ is agent $i$’s $\alpha_{i}$-ideal utility probability function with value distribution $\mathcal{F}_{i}$. (We define $\alpha_{i}$-ideal utility probability function in Definition˜B.1; informally, it is $1$ if $V_{i}[t]$ is in the top $\alpha_{i}$-quantile of $\mathcal{F}_{i}$ and 0 otherwise.) Then, agent $i$ will bid if and only if $\hat{\pi}_{i}$ would bid with the Bernoulli value $\hat{V}_{i}[t]$. In other words, the policy $\pi_{i}$ is simply following the policy $\hat{\pi}_{i}$ but with the Bernoulli values $\hat{V}_{i}[t]$ instead of the actual values.
@@
Therefore, $M[t]$ is a $\mathcal{G}_{t}$-martingale. Let $\bar{v}$ be an upper bound on the distribution $\mathcal{F}_{i}$ (recall we assumed value distributions are bounded). By the Azuma-Hoeffding inequality, for any $\varepsilon>0$,
$\Pr(M[T]\leq-\varepsilon)\leq\exp\left(-\frac{2\varepsilon^{2}}{\bar{v}^{2}T}\right).$
Set $\varepsilon=\bar{v}\sqrt{T\ln T}$, so that the above is at most $O(1/T^{2})$. Let $E_{2}$ be the event that the above does not occur and let $E=E_{1}\cap E_{2}$. We have $\Pr(E)\geq 1-O(1/T^{2})$. On $E$,
@@
thereby proving the lemma.
Now we can prove Lemma˜5.1.
Proof 22 (Proof of Lemma˜5.1).
@@
This will be at most $\alpha_{i}\bar{p}(1-\alpha_{i})T+O(\sqrt{T\log T})$ on an event $E_{1}$ of probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by Azuma-Hoeffding applied to $M_{i}^{T}$. Also, agent $i$ will bid at least $\alpha_{i}T-(\sqrt{T\log T})$ times on some event $E_{2}$ or probability at least $1-O\left(\frac{1}{T^{2}}\right)$ by standard Chernoff bounds. Let $E=E_{1}\cap E_{2}$. We lower bound agent $i$’s utility on $E$. Appealing to Lemma˜E.6, we assume without loss of generality that agent $i$ has a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution. Then, the $\alpha_{i}$-aggressive strategy is simply the strategy of bidding when $V_{i}[t]=1$ if there is budget left. Agent $i$’s utility can then be bounded below on $E$ as
@@
Because the $\alpha_{i}$-ideal utility under a $\mathrm{Bernoulli}(\alpha_{i})$ value distribution is $\alpha_{i}$, this establishes the lemma.
E.6 Completion of the Proof of Lemma˜5.3
@@
where the inequality comes from the inequalities $2+2\prod_{j\notin I}(1-y_{j})\leq 0$ and $1+X\leq 0$.
@@
(E.8) $\begin{split}(y_{j_{1}}-y_{j_{2}})\left(\frac{\partial f}{\partial y_{j_{1}}}-\frac{\partial f}{\partial y_{j_{2}}}\right)=\frac{1}{2}\prod_{i\in I}(1-x_{i})\prod_{j\notin I\cup\{j_{1},j_{2}\}}(1-y_{j})\left(2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\right).\end{split}$
By Lemma˜E.7 applied to the function $x\mapsto\frac{x}{1-x}$ on $[0,1]$,
$2X-\sum_{i\in I}\frac{x_{i}}{1-x_{i}}(1-X)\geq 2X-\frac{\sum_{i\in I}x_{i}}{1-\sum_{i\in I}x_{i}}(1-X)=2X-\frac{X}{1-X}(1-X)=X\geq 0.$
Therefore, (E.8) is nonnegative so $f((x_{i})_{i\in I},\cdot)$ is Schur-convex.
E.7 Deferred Proofs of Propositions˜5.4 and 5.5
Notice that Proposition˜5.4 is a special case of Lemma˜5.5, so we only need to prove Lemma˜5.5.
Proof 24 (Proof of Lemma˜5.5).
@@
For any single agent $j$, their bids are i.i.d. $\mathrm{Bernoulli}(\alpha_{j})$ across time conditioned on them having budget remaining. Let $E$ be the event that no agent $j\neq i$ runs out of budget. For the same reason as in Lemma˜E.2, the probability that agent $j$ runs out of budget is at most $O(1/T^{2})$, so $\Pr(E)\geq 1-O(1/T^{2})$. Using the idea and notation of the imaginary game where budgets are not enforced as introduced in subsection˜E.3, agent $i$’s expected utility can be bounded as
@@
By the strategy of the other agents, (we use the notation for $\beta$-ideal utility as in Definition˜B.1)
@@
using Lemma˜E.1 for the first inequality and Lemma˜E.4 for the second. The above and (24) imply the lemma statement.
E.8 Deferred Proof of Theorem˜5.6
Proof 25 (Proof of Theorem˜5.6).
@@
Using the above and Lemma˜5.5,